A calibrated pipet (95 drops/mL) delivers 12 drops of stearic acid solution. (1.5x10^-4 g/mL) to give a monolayer with a diameter of 12.5 cm. Calculate the surface area of tge monolayer, the moles of stearic acid (284 g/mol) in the monolayer, the number of stearic acid molecules in the monolayer, and the apparent volume & cross-sectional area occupied by each molecule.
Answer #1

Condition 1:

Since the pipette is calibrated to give 95 drops/mL, therefore, in 12 drops there is (12/95 = 0.126 mL = 1.26 x 10^-4 L) of Stearic acid solution (1.5 x 10^-4 g/mL = ((1.5 x 10^-4)/284) mol/L= 5.28 x 10^-7 mol/L).

Condition 2:

Diameter of monolayer = 12.5 cm, Radius r = 6.25 cm.

Answer 1:

Surface area of Momolayer = r2 =  122.71846 cm2.

Answer 2:

Number of moles n = ((1.26 x 10^-4 L ) x (5.28 x 10^-7 mol/L))/1L = 6.65 x 10^-11 moles of stearic acid

Answer 3:

The number of stearic acid molecules in the monolayer = (6.022 x 10^23) x (6.65 x 10^-11) = 4.01 x 10^13 molecules.

Answer 4:

The apparent volume & cross-sectional area occupied by each molecule = (122.71846 cm2)/ (4.01 x 10^13) = 3.06 x 10^-12 cm2